Show that x2 is not O(x*log(x))


1
Expert's answer
2021-10-25T03:25:06-0400

"f(x)=O(g(x))" if there exists a positive real number M and a real number x0 such that

"|f(x)|\le Cg(x)" for all "x\ge x_0"


"|x^2|>xlogx" for positive integers.

So. x2 is not O(x*log(x)).